OpenStudy (anonymous):

How would I solve fraction 2/7 = z/-5?

OpenStudy (unklerhaukus):

you would multiply both sides by negative five

OpenStudy (anonymous):

Or you can transpose: z = 2/7 * -5 = -10/7